OpenStudy (anonymous):

find the dertivative of y in respect to the appropriate variable of y=arcsin(2t)^1/2)

OpenStudy (anonymous):

\[d(\arcsin(t))/dt = 1 / (\sqrt{1 - t ^{2}})\]

zepdrix (zepdrix):

Hey Kaleigh! \(\Large \color{royalblue}{\text{Welcome to OpenStudy! :)}}\) So this is your problem?\[\Large y\quad=\quad \arcsin \sqrt{2t}\]Or this?\[\Large y\quad=\quad \sqrt{\arcsin(2t)}\]

OpenStudy (dan815):

how do you do to the power of more than 1 digit at a time in latex

zepdrix (zepdrix):

Gotta use these fancy curly braces: {} Example \sqrt{2t} --> \(\Large \sqrt{2t}\)